OpenStudy (anonymous):

find the distance between (0,0) and (-2,-2).

OpenStudy (saifoo.khan):

1 - GO to Google .. 2 - Type : ru/mov0001.swf 3 - Click on the first website .. 4 - Enjoy ..

OpenStudy (saifoo.khan):

Satelite, this one is for u.

OpenStudy (anonymous):

vertical distance is 2, and so is horizontal distance. so pythatgoras tells you that \[d^2=2^2+2^2=8\] and your distance is \[\sqrt{8}=2\sqrt{2}\] good morning saifooooooooooooooo

OpenStudy (saifoo.khan):

Good Morning there! xD

OpenStudy (saifoo.khan):

Well, its evening here.

OpenStudy (anonymous):

thanks. i am awake now!

OpenStudy (saifoo.khan):

HAHA!

OpenStudy (saifoo.khan):

How was it?

OpenStudy (anonymous):

d=sqrt[(-2)^2+(-2)^2] =sqrt(8) =2sqrt2

OpenStudy (anonymous):

how do i round that to the neareth tenth